IF THE TIME FOR 50 OSSICILATIONS OF THE PENDULUM IS 80 sec USING A WRIST WATCH OF 1 sec RESOLUTION , THEN ACCURACY IN MEASUREMENT OF TIME PERIOD OF OSCILLATION IS WHAT ?

Grade:12

1 Answers

Vikas TU
4 years ago
% error inmeasurement of time period is :
Error = (Del t / T )*100
Del t = Resolution of watch / No. of oscillation 
Del t = 1/50
T = Total time / No. of oscillation 
T = 80/50
Error = 1/50 * 50/80  *100
Error = 1.25%